OpenStudy (anonymous):

I need help quick. What is the quotient of three and one over five divided by seven over ten in simplest form? three and two over seven three and seven over fifty four and two over five four and four over seven

OpenStudy (mathlegend):

Wait all of those together are being divided?

OpenStudy (anonymous):

What do you mean by 3 and 1? 31?

OpenStudy (anonymous):

please use equation instead of words. hard to analyze ..

OpenStudy (anonymous):

\[3\frac{ 1 }{ 5 }\]

OpenStudy (mathlegend):

okay

OpenStudy (anonymous):

ahh.. much better

OpenStudy (mathlegend):

If I were you I would change that Mixed Number into an improper fraction

OpenStudy (mathlegend):

So can you tell me what \[3\frac{ 1 }{ 5}\] is as an improper fraction?

OpenStudy (anonymous):

no?

OpenStudy (anonymous):

Multiply 3 by 5 and add 1 to get the numerator.

OpenStudy (anonymous):

16

OpenStudy (mathlegend):

Don't forget to include the denominator.

OpenStudy (anonymous):

Yep so its 16/5

OpenStudy (anonymous):

as an improper fraction.

OpenStudy (mathlegend):

\[\frac{ 16 }{ 5}\]*\[\frac{ 10 }{ 7}\]

OpenStudy (anonymous):

you will make the \[3\frac{ 1 }{ 5 }\] in a improper fraction then switch the numerator of the second fraction then multiply it

OpenStudy (mathlegend):

When dividing I like to flip the second fraction (Known as the reciprical)

OpenStudy (mathlegend):

This way you can multiply

OpenStudy (anonymous):

correction: reciprocal

OpenStudy (mathlegend):

16/5 * 10/7 = ?

OpenStudy (mathlegend):

Multiply Numerators and then Multiply the denominators

OpenStudy (anonymous):

160 35

OpenStudy (anonymous):

What @MathLegend says is correct. \[\frac{ 16 }{ 5 } \times \frac{ 10 }{ 7 } = \frac{ 160 }{ 35 }\]

OpenStudy (mathlegend):

Good

OpenStudy (anonymous):

Can you simplify on your own now?

OpenStudy (anonymous):

instead of using \[\frac{ 16 }{ 5 }\] times \[\frac{ 10 }{ 7 }\] = \[\frac{ 160 }{ 35 }\] .. use the cross multiplication for easy answering.. it will help you ..

OpenStudy (mathlegend):

It is the same thing... doesn't matter.